OpenStudy (anonymous):

how do u solve 9a (squared) - 30 =3a

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

first you can see that you can divide it by 3 and put everything to the left hand side to get 3a^2-a-10=0

OpenStudy (anonymous):

now you can factor it or just use the formula for quadratic expressions

OpenStudy (amistre64):

complete squares is easiest

OpenStudy (amistre64):

\(9a^2 - 30 =3a ;\) +30, -3a, /9 \[9(a^2 -{1\over3}a) = 30 \] \(9(a-{1\over6})^2 = 30 +{9\over36}\); /9 \[a-{1\over6} = \pm \sqrt{{10\over3} +{1\over36}}\] \[a = {1\over6} \pm {\sqrt{13}\over6}\] maybe lol

OpenStudy (amistre64):

the latex slows me down